OpenStudy (anonymous):

.Input the equation of the given line in slope-intercept form. The line with m = 2 and b = -4.

OpenStudy (anonymous):

y = mx+b put m and b

OpenStudy (anonymous):

y=2x-4?

OpenStudy (anonymous):

or y=2x+1?

OpenStudy (anonymous):

First option

OpenStudy (anonymous):

what about this problem? Input the equation of the given line in standard form. The line with m = 2/3 and passing through (1, 1)

OpenStudy (anonymous):

\[(y-y_1) = m(x-x_1)\]

OpenStudy (anonymous):

same formula for this one to? Input the equation of the given line in standard form. The line through (1, -4) and parallel to 2x + 3y = 4.

OpenStudy (anonymous):

yes use above formula. and for m, parallel lines have same m

OpenStudy (anonymous):

the first answer would be just x right?

OpenStudy (anonymous):

\[(y-y_1)=m(x-x_1)\] for m, parallel lines have same m

OpenStudy (anonymous):

i dont understand how to do the problem still ;(

OpenStudy (anonymous):

The line through (1, -4) and parallel to 2x + 3y = 4. line parallel to 2x + 3y = 4. will have same slope as of the given line to find slope of the line. put the eq. in y=mx+c form

OpenStudy (anonymous):

i got the first problem now I just need the answer to Input the equation of the given line in standard form. The line through (1, -4) and parallel to 2x + 3y = 4.

OpenStudy (anonymous):

The line through (1, -4) and parallel to 2x + 3y = 4. line parallel to 2x + 3y = 4. will have same slope as of the given line to find slope of the line. put the eq. in y=mx+c form y =-2/3x +4/3 so the slope is -2/3 slope of new line will also be -2/3 \[(y-y_1) = m(x-x_1)\] m = -2/3 (x1,y1) = (1,-4)